solve differential equation dy/dx=(2x^2+y^2)/(-2xy+3y^2)
1
Expert's answer
2021-08-04T12:32:28-0400

Let us solve the differential equation "\frac{dy}{dx}=\frac{2x^2+y^2}{-2xy+3y^2}," which is equivalent to "\frac{dy}{dx}=\frac{2+(\frac{y}{x})^2}{-2\frac{y}{x}+3(\frac{y}{x})^2}." Let us use the transformation "y=ux." Then "\frac{dy}{dx}=\frac{du}{dx}x+u." We get the equation "\frac{du}{dx}x+u=\frac{2+u^2}{-2u+3u^2}," which is equivalent to "\frac{du}{dx}x=\frac{2+u^2}{-2u+3u^2}-u," and hence to "\frac{du}{dx}x=\frac{2+u^2+2u^2-3u^3}{-2u+3u^2}." Then we have the equation "\frac{-2u+3u^2}{2+3u^2-3u^3}du=\frac{dx}{x}." It follows that


"\int\frac{-2u+3u^2}{2+3u^2-3u^3}du=\int\frac{dx}{x}"


"-\frac{1}{3}\int\frac{6u-9u^2}{2+3u^2-3u^3}du=\int\frac{dx}{x}"


"-\frac{1}{3}\int\frac{d(2+3u^2-3u^3)}{2+3u^2-3u^3}=\int\frac{dx}{x}"


"\ln|x|=-\frac{1}{3}\ln|2+3u^2-3u^3|+\ln|C|"


"\ln|x|+\frac{1}{3}\ln|2+3u^2-3u^3|=\ln|C|"


"\ln|x(2+3u^2-3u^3)^{\frac{1}{3}}|=\ln|C|"


"x\sqrt[3]{2+3u^2-3u^3}=C"


"x\sqrt[3]{2+3(\frac{y}{x})^2-3(\frac{y}{x})^3}=C"


"\sqrt[3]{2x^3+3y^2x-3y^3}=C"


It follows that the general solution of the differential equation "\frac{dy}{dx}=\frac{2x^2+y^2}{-2xy+3y^2}" is


"2x^3+3y^2x-3y^3=C."